# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\underline{\rm mrid}_{\mu}(f,K,d)$ for every $\mu\in K$.
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then, by Theorem 1.2, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. This shows that
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
@@
$\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
This completes the proof of (3.18).
Next we show the second equality:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
Using formula (3.3), the inequality $\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is clear. On the other hand, let $\gamma>0$ be sufficiently small and put
$M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$
Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
@@
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows that
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}.$
Letting $\gamma\to 0$, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
@@ 3.3 Upper capacity metric mean dimension of saturated sets
In this subsection, we prove the Theorem 1.4.
By Proposition 2.1, the inequality
@@
In light of Theorem 2.4, we need to show that for $\epsilon^{\ast}>0$, $\zeta>0$ sufficiently small and $\delta\in(0,1)$, one has
@@
However, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
@@
#$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
@@
For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
@@
Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain the fractal set
@@
It is worth mentioning that for any $(k,j)>(0,0)$, #$\#\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
@@
Hence, for sufficiently large $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
@@
This implies the inequality (3.19).
3.4 An extension to metric mean dimensions of sup sets and sub sets
In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets. The main results are presented in Theorems 3.7, 3.8 and 3.9.
In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the sense of weak-topology. The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01]. of $f$. By Birkhoff ergodic theorem, although the historic set is usually regarded as negligible in the measure-theoretic context, its topological dynamics can be quite rich. Especially for systems with specification-like properties, it can even have full topological entropy [CTS05] and full metric mean dimension [LV21, LL24]. To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03]. More precisely, given a nonempty subset $C\subset M_{f}(X)$, we define the sup set, equ set [^2] In this context, we do not take into account the deformation of $\mathcal{E}_{n}(x)$ via a continuous affine map from $M_{f}(X)$ to a topological vector space $Y$. Hence, we prefer to use the term “saturated set” rather than “equ set”. and sub set of $C$ as
${}^{C}G:$ =$=\{x\in X:V_{f}(x)\supset C\},$
$G_{C}:$ =$=\{x\in X:V_{f}(x)=C\},$
$G^{C}:$ =$=\{x\in X:V_{f}(x)\subset C\},$
respectively.
Note that the equ set is precisely the saturated set associated with $C$ defined previously. More generally, by using two nonempty subsets $C_{1},C_{2}\subset M_{f}(X)$, we define the set
@@ 3.4 An extension to metric mean dimensions of sup sets and sub sets
which covers the three different notions above in an unified manner.
Now, we are ready to establish variational principles for the three given types of level sets $G^{C},^{C}G,G(C_{1},C_{2})$ by means of the Bowen metric mean dimension.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty subset $C\subseteq M_{f}(X)$,
@@
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
where $\overline{{\rm Conv}}(C)$ denotes the convex closure of set $C$.
Indeed, since $C\subset\overline{{\rm Conv}}(C)$, it suffices to show
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Let $\mu\in\overline{{\rm Conv}}(C)$. We can choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $\lambda_{n}^{(j)}>0$ and $\sum_{j=o}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$. The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf. [FW16, Lemma 2.3, (3)]) implies that
@@
Combining this and $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is concave in $\mu$, we can deduce that
$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma$
holds for any for $\gamma>0$. This yields the desired inequality and hence (3.20) holds.
Notice that $G_{\overline{{\rm Conv}}(C)}\subset^{C}G$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C)},d\right)$
@@
Conversely, for any $\mu\in C$ and $\epsilon>0$, by Lemma 3.6 one has
$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This shows $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ for every $\epsilon>0$. ∎
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty closed subset $C\subseteq M_{f}(X)$, we have
@@
For every $\epsilon>0$, by Lemma 3.6 again we have
@@
By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon)$
holds for all $\mu\in C$ and any sufficiently small $\epsilon>0$. Combining with the fact that $G_{\mu}\subset G^{C}$ holds for each $\mu\in C$, one gets
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Suppose that $C_{1}\subseteq M_{f}(X)$ is a nonempty set satisfying that $\overline{{\rm Conv}}(C_{1})$ is a connected component of $C_{2}\subset M_{f}(X)$. Then
@@
Since $C_{1}\subset\overline{{\rm Conv}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C_{1})},d\right)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm Conv}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
@@
Let $(X,d,f)$ be a t.d.s. with specification property.
@@
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
@@
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
@@
(1) follows by Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
@@
Let $(X,d,f)$ be a t.d.s. with specification property.
@@
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
@@
    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
@@
    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
@@
$(1)$ is clear. For $(2)$, it follows from Theorem 1.1 that
@@
$(3)$ follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
4 Some applications of main results
In this section, we apply the main results to multifractal analysis of level sets and mean Li-Yorke chaos in infinite entropy systems. The main results in this section are presented in Theorems 4.1 and 4.3.
4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ within the context of an abstract framework.
Given a dynamical system $(X,f)$, assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function. We define
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right].$
For any $\alpha\in L_{\varphi}$, we define $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$. The continuity of $\varphi$ then implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$. We define the level set of $\alpha$ w.r.t. $\varphi$ as
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\},$
which consists of the set of points of $X$ such that $\varphi$ takes constant value $\alpha$ on $V_{f}(x)$.
Condition A: For every $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Several examples of the function $\varphi$ that satisfy Condition A are stated below:
  • (a)
    If $\varphi=\alpha\in\mathbb{R}$ is a constant, then $L_{\varphi}=\{\alpha\}$, $K_{\alpha}=X$, $M_{f}(\varphi,\alpha)=M_{f}(X)$;
  • (b)
    We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu},$
    where $\psi$, $\phi\in C(X)$ and $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$).
    Then for any $\alpha\in L_{\varphi}$,
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\},$
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\};$
  • (c)
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, that is, for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
    $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for all $\alpha\in L_{\Phi}$. Specially, let $\psi\in C(X)$ and define $\phi_{n}=S_{n}\psi$, then $\Phi=\{\phi_{n}\}$ is an additive potential, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given a nonempty open subset $U\subset X$, we similarly define the mixed $\alpha$-level set
$K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$

### v4

@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d)$.
To show the reverse inequality
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\geq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
we assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. By Theorem 1.2, we can choose a strictly decreasing sequence $\epsilon_{k}\to 0$ as $k\to 0$ and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$ (3.18)
Without loss of generality, we assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. For every $\epsilon>0$, if $\epsilon\in(\epsilon_{k+1},\epsilon_{k}]$ for some $k\geq 1$, we define $\nu_{\epsilon}:=\epsilon_{k}$. Then $(\nu_{\epsilon})\in M_{K}(\mu)$. Using (3.18), we have
$\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
@@
$\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
Next we show the variational principle:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\limits\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
Using (3.3), there exists a constant $c>0$ for every sufficiently small $\epsilon>0$,
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K},c\epsilon).$
This yields the inequality
$\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
On the other hand, let $\gamma>0$ and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$ By Theorem 1.1, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real numbers that converges to 0 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
@@
By the similar arguments, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
@@ 3.3 Upper capacity metric mean dimension of saturated sets
In this subsection, we give the proof of Theorem 1.4.
By Proposition 2.1, it is clear that
@@
In light of Theorem 2.4, we need to show that for every sufficiently small $\epsilon^{\ast}>0$, $\zeta>0$ and $\delta\in(0,1)$, one has
@@
Here, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
@@
$S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
@@
For the non-empty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with ${\overline{B}(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$, and $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
@@
Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain a fractal set
@@
It is worth mentioning that for any $(k,j)>(0,0)$, $\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
@@
Hence, $\mathbb{F}$ is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset with $S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus,
@@
This implies the inequality (3.19) by letting $\zeta\to 0$.
3.4 Metric mean dimensions of sup sets and sub sets
In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets.
In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the weak-topology. The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01]. of $f$. To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03]. More precisely, given a non-empty subset $C\subset M_{f}(X)$, we define the sup set and sub set of $C$ as
$^{C}G:$ $=\{x\in X:V_{f}(x)\supset C\},$
$G^{C}:$ $=\{x\in X:V_{f}(x)\subset C\},$
respectively. More generally, for two non-empty subsets $C_{1},C_{2}\subset M_{f}(X)$, we can define
@@ 3.4 Metric mean dimensions of sup sets and sub sets
to recover the above notions in an unified manner.
Now, we are ready to establish variational principles for $G^{C},^{C}G$.
@@
Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty subset $C\subseteq M_{f}(X)$,
@@
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
where $\overline{{\rm co}}(C)$ denotes the closure of the convex hull of $C$.
Indeed, since $C\subset\overline{{\rm co}}(C)$, it suffices to show
$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Let $\mu\in\overline{{\rm co}}(C)$. Choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $0<\lambda_{n}^{(j)}<1$ and $\sum_{j=0}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$. The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf. [FW16, Lemma 2.3, (3)]) implies that
@@
Since $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is a concave function on $M_{f}(X)$, for every $\gamma>0$ we can deduce that
$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma.$
This yields the desired inequality (3.20).
Notice that $G_{\overline{{\rm co}}(C)}\subset^{C}G$, we have
$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C)},d\right)$
@@
Fix $\mu\in C$ and $\epsilon>0$. By Lemma 3.6, one has
$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
and hence $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
@@
Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty closed subset $C\subseteq M_{f}(X)$, we have
@@
For every $\epsilon>0$, by Lemma 3.6 we have
@@
By (3.17), there exists a constant $c>0$ such that for all $\mu\in C$ and any sufficiently small $\epsilon>0$,
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
Combining this inequality with the fact that $G_{\mu}\subset G^{C}$ for each $\mu\in C$, we have
@@
Let $(X,d,f)$ be a TDS with the specification property. Let $C_{1}\subseteq M_{f}(X)$ be a non-empty set and $C_{1}\subseteq M_{f}(X)$. If $\overline{{\rm co}}(C_{1})$ is a connected component of $C_{2}$, then
@@
Since $C_{1}\subset\overline{{\rm co}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C_{1})},d\right)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm co}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
@@
Let $(X,d,f)$ be a TDS with the specification property.
@@
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
@@
    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
@@
(1) follows from Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
@@
Let $(X,d,f)$ be a TDS with the specification property.
@@
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
@@
    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
@@
    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
@@
$(1)$ is clear.
$(2)$. It follows from Theorem 1.1 that
@@
$(3)$. It follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
4 The applications of main results
In this section, we apply the main results to study the metric mean dimension of level sets, mean Li-Yorke chaos in infinite entropy systems and the metric mean dimension of the set of generic points of full shifts over compact metric spaces.
4.1 Metric mean dimensions of level sets
Let $(X,d,f)$ be a TDS and $\psi$ be a continuous function on $X$. If $\mu$ is a $f$-invariant probability measure on $X$, the Birkhoff ergodic theorem verifies that for $\mu$-a.e. $x\in X$, the limit
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)$
exists.
However, given $\alpha\in\mathbb{R}$ it does not give more information about the “size” of the set of the points of $X$ that the time average of $\psi$ equals $\alpha$. For this reason, the $\alpha$-level set of $\psi$ is defined by
$K_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}.$
In [BR23], the authors proved that if $K_{\alpha}$ is non-empty and the TDS $(X,d,f)$ has the specification property, then
$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $M_{f}(\psi,\alpha):=\{\mu\in M_{f}(X):\int\psi d\mu=\alpha\}$.
Actually, we are able to prove this variational principle by an alternative approach and give new characterization for the (packing/upper capacity) metric mean dimension of level sets.